# Changes in Real Time: Online Scene Change Detection with Multi-View Fusion

### v2

@@ Abstract
Online Scene Change Detection (SCD) is an extremely challenging problem that requires an agent to detect relevant changes on the fly while observing the scene from unconstrained viewpoints. Existing online SCD methods are significantly less accurate than offline approaches. We present the first online SCD approach that is pose-agnostic, label-free, and ensures multi-view consistency, while operating at over 10 FPS and achieving new state-of-the-art performance, surpassing even the best offline approaches. Our method introduces a new self-supervised fusion loss to infer scene changes from multiple cues and observations, PnP-based fast pose estimation against the reference scene, and a fast change-guided update strategy for the 3D Gaussian Splatting scene representation. Extensive experiments on complex real-world datasets demonstrate that our approach outperforms both online and offline baselines. Our code will be released at O-SCD.github.io.
@@ 1 Introduction
Detecting changes in a scene is an essential task in scene understanding, with numerous applications in environmental monitoring [taneja2011image], infrastructure inspection [han2021change], and damage assessment [sakurada_change_2015]. Scene change detection (SCD) is especially challenging in the context of robotics, where an agent observes the scene from unconstrained and independent viewpoints when re-visiting it after some time, while having to discern relevant (e.g. object movement) and irrelevant changes (e.g. caused by shadows, or reflections).
@@
To address these challenges, recent approaches [galappaththige2025multi, lu20253dgs, jiang2025gaussian, zhou2023pad, kruse2024splatpose] leverage photorealistic 3D scene representations like Neural Radiance Fields [mildenhall2021nerf] (NeRF) and 3D Gaussian Splatting [kerbl20233d] (3DGS) to enable pose-agnostic SCD from unconstrained viewpoints. Complementary efforts [galappaththige2025multi, kim2025towards, cho2025zero] explore label-free SCD to remove the reliance on costly and labor-intensive human-labeled changes, improving robustness under domain and data distribution shifts.
Despite these advances, state-of-the-art (SOTA) SCD methods [kim2025towards, lu20253dgs, galappaththige2025multi] are limited to an offline setting, where both pre- and post-change scene captures are available in advance. Online methods [liu2024splatpose+, kruse2024splatpose, park2021changesim, park2021changesim] can detect changes on the fly, as new images are acquired while the agent revisits the scene, enabling the use of SCD for real-time decision making and intervention. As illustrated in Fig. 1, the change detection performance of these online approaches remains substantially lower than SOTA offline methods, while many fail to sustain real-time performance.
We introduce a novel SCD approach that, for the first time, unifies the strengths of online, pose-agnostic, and label-free methods, while additionally enforcing multi-view consistency [galappaththige2025multi]. Our method surpasses existing online and offline methods in detection performance while operating at over 10 FPS. This leap in both speed and accuracy is enabled by two key innovations: a novel self-supervised loss enforcing change information consistency across viewpoints, addressing the limitations of hard-thresholded intersection fusion [galappaththige2025multi] (see Fig. 2), and an ultra-light PnP-based pose estimation module.
@@
112233445566778899101011111212131300.20.20.40.40.60.6MV3DCD [galappaththige2025multi] (Offline)GeSCD [kim2025towards] (Offline)ChangeSim (CS)[park2021changesim]CS+CYWS2D[cyws2d]CS+GeSCD[kim2025towards]OmniPoseAD[zhou2023pad]SplatPose[kruse2024splatpose]SplatPose+[liu2024splatpose+]OursFrames per Second (FPS)Change Detection Performance(F1 Score)
@@
Figure 1: Our online Scene Change Detection method establishes a new state of the art, detecting changes more reliably than all prior methods, including the strongest offline baselines. It operates at a runtime comparable to the fastest online approaches while achieving substantially higher F1 scores. These gains are enabled by a self-supervised loss enforcing multi-view consistency and a lightweight PnP-based pose estimation module.
@@
Figure 2: Qualitative comparison with MV3DCD [galappaththige2025multi]. MV3DCD’s hard thresholding and intersection heuristic lead to missed or spurious detections, especially for subtle appearance changes in semantically similar objects (red-to-blue T-shaped object in Meeting Room, blue-to-black bench in Porch). Hard thresholding risks discarding subtle but important changes, while the intersection fails to capture true changes unless present in both masks. Our method jointly learns complementary change information in pixel- and feature-level cues via our novel self-supervised loss, capturing fine-grained changes and achieving state-of-the-art performance in both online and offline settings.
@@
Our third innovation is a change-guided update strategy for the 3DGS-based scene representation. Maintaining an up-to-date representation of an evolving scene without naively reconstructing it from scratch is challenging [ackermann2025clsplats, cai2023clnerf, zeng2025gaussianupdate, cheng2025lt], but essential for long-term monitoring. Naïve reconstruction after each inspection round is computationally expensive and discards well-reconstructed information from unchanged regions. We address this by leveraging the predicted change masks to guide selective updates: only changed regions are newly reconstructed, fused with existing primitives, and refined through a lightweight global adjustment while preserving the geometry and the appearance of unchanged areas. Our selective update approach enables scene representation updates in seconds while reusing the high-fidelity representations of unchanged regions.
@@
In summary, we make three contributions validated by our extensive experiments across indoor and outdoor real-world environments:
@@ 2.1 Scene Change Detection
SCD has traditionally been studied as a bitemporal pairwise problem, where a model detects changes between two images captured at two time instances from identical [alcantarilla2018street, daudt2018fully, sakurada_change_2015, chen2021dr, varghese2018changenet, lei2020hierarchical] or closely-aligned [cyws2d, cyws3d, lin2025robust] viewpoints. Most approaches formulate this as a segmentation task [alcantarilla2018street, daudt2018fully, sakurada_change_2015, lin2025robust, chen2021dr], while few explore bounding box prediction [cyws2d, cyws3d], relying fully [alcantarilla2018street, daudt2018fully, sakurada_change_2015, chen2021dr, varghese2018changenet, lei2020hierarchical, cyws2d, cyws3d, lin2025robust] or partially [lee2024semi, sakurada2020weakly] on costly human annotations to compensate for lighting variations, seasonal changes, or viewpoint inconsistencies. However, this paradigm has clear limitations: performance degrades under distribution shifts, annotation is tedious, and the range of possible changes in complex scenes is virtually unbounded. Recent work [galappaththige2025multi, kim2025towards, cho2025zero, kannanZero, alpherts2025emplace] has shifted toward label-free or zero-shot approaches, driven by the emergence of powerful visual foundation models [kirillov2023segment, oquab2023dinov2]. However, these methods [kim2025towards, cho2025zero, alpherts2025emplace, kannanZero] assume image pairs with identical viewpoints are available—a condition rarely satisfied in autonomous systems operating along independent trajectories, and a condition we do not assume.
@@
More recent methods [galappaththige2025multi, lu20253dgs, jiang2025gaussian] exploit high-fidelity 3D scene representations [kerbl20233d, wu20244d] to model the pre-change scene and render novel views from post-change viewpoints, motivating pose-agnostic SCD. MV3DCD [galappaththige2025multi] showed that learning change information across multiple viewpoints with a scene representation significantly outperforms pairwise predictions. However, these approaches require a complete set of pre- and post-change captures, and use Structure-from-Motion [schonberger2016structure] (SfM) to register poses to a common reference frame, confining them to an offline setting. In contrast, we study SCD in an online and incremental regime, where changes are inferred on-the-fly.
@@
Pose-agnostic anomaly detection [zhou2023pad, kruse2024splatpose, liu2024splatpose+] also builds a 3D representation of a pre-change object. To detect anomalies, the object is rendered from the post-change viewpoint, then scored using feature comparisons. However, these works focus on single objects rather than large-scale scenes. Approaches such as [kruse2024splatpose, zhou2023pad] optimize camera poses directly against the representation, leading to slower pose estimation and, as shown by MV3DCD [galappaththige2025multi], frequent convergence failures in large complex scenes with multiple changes and view-dependent inconsistencies. Liu et al[liu2024splatpose+] improved efficiency by replacing this step with HLoc [sarlin2019coarse].
@@
ChangeSim [park2021changesim] also investigates online SCD. However, ChangeSim depends on an off-the-shelf RGB-D SLAM system [labbe2019rtab] for pose estimation and assumes that pre- and post-change trajectories are closely aligned. This reduces the task to image retrieval, where the nearest pre-change view (by $L_{1}$ distance between camera poses) is selected, followed by pairwise prediction using an existing supervised SCD model [sakurada2020weakly, varghese2018changenet]. In contrast, we make no assumptions about incoming RGB-only frames or the trajectories; instead, we estimate poses directly in the pre-change coordinate frame and infer change masks by jointly leveraging all viewpoints observed so far. Our approach operates fully label-free, pose-agnostic, multi-view, online, and at real-time rates.
@@ 2.2 Efficient Representation Update
NeRFs [mildenhall2021nerf] and 3DGS [kerbl20233d] are widely adopted photorealistic scene representations, capturing fine geometry and appearance. NeRFs regress a 5D plenoptic function [bergen1991plenoptic] using an MLP network to parameterize density and view-dependent radiance, while 3DGS employs anisotropic Gaussian primitives for real-time novel-view-synthesis.
Recently, there has been growing interest in real-time reconstruction [meuleman2025fly, lin2025longsplat]. However, these methods generally underperform compared to offline counterparts and require substantial view overlap between frames. In contrast, approaches that address the long-term evolution of scenes remain less explored, focusing on updating representations from sparse and intermittent captures [ackermann2025clsplats, wu2023cl]. Cheng et al[cheng2025lt] proposed a multimodal strategy incorporating LiDAR to enable targeted updates in 3DGS. Closely related is continual learning for photorealistic scene representations [ackermann2025clsplats, zeng2025gaussianupdate, cai2023clnerf, wu2023cl]. CL-NeRF [wu2023cl] introduces lightweight expert adaptors with conflict-aware distillation, while CLNeRF [cai2023clnerf] adopts generative replay with incremental updates in NeRFs. However, these approaches inherit NeRF’s slow inference and long runtimes. GaussianUpdate [zeng2025gaussianupdate] proposes a three-stage optimization pipeline requiring substantial training iterations. CL-Splats [ackermann2025clsplats] introduces a local optimization kernel to calculate gradients only for changed primitives, yet cannot robustly handle global appearance variations, particularly the illumination shifts often present between real-world inspection scenarios.
Continual learning approaches [ackermann2025clsplats, zeng2025gaussianupdate, wu2023cl, cai2023clnerf] focus on updating representations while facilitating history recovery. In contrast, we focus on updating the representation with minimal training overhead, enabling frequent repeated inspections. To this end, we propose a simple selective modeling strategy that only reconstructs changed regions, guided by our change masks and the change representation, followed by fusion with existing primitives. A lightweight global optimization step ensures consistency, enabling updates within seconds while robustly handling both geometric and appearance changes, including global illumination variations.
@@ 3 Methodology
Our approach is illustrated in Fig. 3. We begin by constructing a 3DGS [kerbl20233d] representation of the pre-change (reference) scene offline (Sec. 3.1). Incoming images of the post-change (inference scene) are processed online. We estimate its pose relative to the reference scene (Sec. 3.2), then render the corresponding viewpoint to extract change cues (Sec. 3.3). These cues are used to infer a change mask (Sec. 3.4), leveraging current and previous observations. After processing all observations, the representation is updated (Sec. 3.5) to the current state of the environment.
Contributions: While our method is based on the SOTA MV3DCD [galappaththige2025multi], we make several important contributions that improve both the runtime and SCD performance. Instead of using MV3DCD’s heuristic fusion method with hard thresholds to enforce multi-view consistency, we introduce a novel self-supervised loss (Sec. 3.4) which improves the SCD mean intersection over union (mIoU) by approximately 15.5% in an offline setting. Furthermore, we replace MV3DCD’s SfM [schonberger2016structure]-based pose estimation (which requires joint access to all pre- and post-change captures) with an ultra-light PnP-based method (Sec. 3.2) which enables online SCD at real-time rates and runs approximately two orders of magnitude faster. Additionally, a new change-guided update strategy (Sec. 3.5) speeds up the representation update by over an order of magnitude compared to optimization from scratch using vanilla 3DGS [kerbl20233d].
Problem Setup: The reference scene $\mathcal{R}_{\text{ref}}$ is captured with a set of $n_{\text{ref}}$ images, $\mathcal{I}_{\text{ref}}=\{I_{\text{ref}}^{k}\}_{k=1}^{n_{\text{ref}}}$. Over time, the scene undergoes changes in structure (e.g., additions, removals, or object movement), or object appearance (e.g., variations in color or texture), forming the inference scene $\mathcal{R}_{\text{inf}}$. In addition, there may be ‘distractors’ such as reflections, shadows, and global illumination changes.
@@ 3.1 Building Reference Scene Representation
Following the standard 3DGS pipeline [kerbl20233d], we first estimate the camera poses $\mathcal{P}_{\text{ref}}=\{P_{\text{ref}}^{k}\}_{k=1}^{n_{\text{ref}}}$ for all images in $\mathcal{I}_{\text{ref}}$ using SfM [schonberger2016structure]. Using $\mathcal{I}_{\text{ref}}$, $\mathcal{P}_{\text{ref}}$ and a sparse point cloud from SfM, we construct the $\mathcal{R}_{\text{ref}}$ with Speedy-Splat [HansonSpeedy]. We assume that the viewpoints, scene coverage, and image quality in $\mathcal{I}_{\text{ref}}$ are sufficient to produce a high-fidelity $\mathcal{R}_{\text{ref}}$.
@@ 3.2 Inferring Pose of an Incoming Frame
For each reference image $I_{\text{ref}}^{k}\in\mathcal{I}_{\text{ref}}$, we extract keypoints and descriptors using XFeat [potje2024xfeat] as a fast, lightweight detector, followed by exhaustive matching across all reference images. Using the known camera poses $\mathcal{P}_{\text{ref}}$, these correspondences are triangulated to form a consistent 3D point set associated with each $I_{\text{ref}}^{k}$. This point set serves as a geometric anchor for estimating the pose of incoming inference frames by establishing 2D–3D correspondences between their detected keypoints and the reference points.
Given an incoming frame $I_{\text{inf}}^{k}$, we extract its descriptors and select the top-$n$ reference frames with the highest number of matches ($n=4$ in our experiments). These reference frames provide candidate 2D–3D correspondences, which are then used to estimate the pose $P_{\text{inf}}^{k}$ of $I_{\text{inf}}^{k}$ via PnP with RANSAC [lepetit2009epnp, fischler1981ransac]. Finally, we refine $P_{\text{inf}}^{k}$ using all inliers through a GPU-parallel mini bundle adjustment [meuleman2025fly].

### v3

@@ Abstract
Online Scene Change Detection (SCD) is an extremely challenging problem that requires an agent to detect relevant changes on the fly while observing the scene from unconstrained viewpoints. Existing online SCD methods are significantly less accurate than offline approaches. We present the first online SCD approach that is pose-agnostic, label-free, and ensures multi-view consistency, while operating at over 10 FPS and achieving new state-of-the-art performance, surpassing even the best offline approaches. Our method introduces a new self-supervised fusion loss to infer scene changes from multiple cues and observations, PnP-based fast pose estimation against the reference scene, and a fast change-guided update strategy for the 3D Gaussian Splatting scene representation. Extensive experiments on complex real-world datasets demonstrate that our approach outperforms both online and offline baselines. Code is available at https://chumsy0725.github.io/O-SCD/.
@@ 1 Introduction
Detecting changes in a scene is an essential task in scene understanding, with numerous applications in environmental monitoring [46], infrastructure inspection [14], and damage assessment [40]. Scene change detection (SCD) is especially challenging in the context of robotics, where an agent observes the scene from unconstrained and independent viewpoints when re-visiting it after some time, while having to discern relevant (e.g. object movement) and irrelevant changes (e.g. caused by shadows, or reflections).
@@
To address these challenges, recent approaches [13, 30, 17, 52, 22] leverage photorealistic 3D scene representations like Neural Radiance Fields [32] (NeRF) and 3D Gaussian Splatting [19] (3DGS) to enable pose-agnostic SCD from unconstrained viewpoints. Complementary efforts [13, 20, 8] explore label-free SCD to remove the reliance on costly and labor-intensive human-labeled changes, improving robustness under domain and data distribution shifts.
Despite recent advances, state-of-the-art (SOTA) SCD methods [20, 30, 13] are confined to an offline setting, where both pre- and post-change observations are available prior to inference. In contrast, online change inference—referred to as online SCD—detects changes on the fly as new images are acquired during a scene revisit, without access to future observations. This setting is critical for real-time decision-making and intervention in embodied and robotic systems. As in prior online approaches [29, 22, 34, 34], we assume that the scene remains static within each revisit. As shown in Fig. 1, existing online SCD methods exhibit a substantial accuracy gap compared to SOTA offline approaches. Moreover, many fail to sustain real-time performance, limiting their practical applicability.
@@
112233445566778899101011111212131300.20.20.40.40.60.6MV3DCD [13] (Offline)GeSCD [20] (Offline)ChangeSim (CS)[34]CS+CYWS2D[39]CS+GeSCD[20]OmniPoseAD[52]SplatPose[22]SplatPose+[29]OursFrames per Second (FPS)Change Detection Performance(F1 Score)
@@
Figure 1: Our online scene change detection method establishes a new state of the art, detecting changes more reliably than all prior methods, including the strongest offline baselines. It operates at a runtime comparable to the fastest online approaches while achieving substantially higher F1 scores.
@@
Figure 2: Qualitative comparison with MV3DCD [13]. MV3DCD’s hard thresholding and intersection heuristic lead to missed or spurious detections, especially for subtle appearance changes in semantically similar objects (red-to-blue T-shaped object in Meeting Room, blue-to-black bench in Porch). Hard thresholding risks discarding subtle but important changes, while the intersection fails to capture true changes unless present in both masks. Our method jointly learns complementary change information in pixel- and feature-level cues via our novel self-supervised loss, capturing fine-grained changes and achieving state-of-the-art performance in both online and offline settings.
We introduce a novel SCD approach that, for the first time, unifies the strengths of online, pose-agnostic, and label-free methods, while additionally enforcing multi-view consistency [13] during inference. Our method surpasses existing online and offline methods in detection performance while operating at over 10 FPS. This leap in both speed and accuracy is enabled by two key algorithmic and system-level innovations: a novel self-supervised loss enforcing change information consistency across viewpoints, addressing the limitations of hard-thresholded intersection fusion [13] (see Fig. 2), and an ultra-light PnP-based pose estimation module.
@@
Our third innovation is a change-guided update strategy for the 3DGS-based scene representation. Maintaining an up-to-date representation of an evolving scene without naively reconstructing it from scratch is challenging [1, 5, 51, 7], but essential for long-term monitoring. Naïve reconstruction after each inspection round is computationally expensive and discards well-reconstructed information from unchanged regions. We address this by leveraging the predicted change masks to guide selective updates: only changed regions are newly reconstructed, fused with existing primitives, and refined through a lightweight global adjustment while preserving the geometry and the appearance of unchanged areas. Our selective update approach enables scene representation updates in seconds while reusing the high-fidelity representations of unchanged regions.
@@
In summary, we make three contributions validated by our extensive experiments across real-world environments:
@@ 2.1 Scene Change Detection
SCD has traditionally been studied as a bitemporal pairwise problem, where a model detects changes between two images captured at two time instances from identical [2, 9, 40, 6, 47, 25] or closely-aligned [39, 38, 28] viewpoints. Most approaches formulate this as a segmentation task [2, 9, 40, 28, 6], while few explore bounding box prediction [39, 38], relying fully [2, 9, 40, 6, 47, 25, 39, 38, 28] or partially [24, 41] on costly human annotations to compensate for lighting variations, seasonal changes, or viewpoint inconsistencies. However, this paradigm has clear limitations: performance degrades under distribution shifts, annotation is tedious, and the range of possible changes in complex scenes is virtually unbounded. Recent work [13, 20, 8, 18, 3] has shifted toward label-free or zero-shot approaches, driven by the emergence of powerful visual foundation models [21, 33]. However, these methods [20, 8, 3, 18] assume image pairs with identical viewpoints are available—a condition rarely satisfied in autonomous systems operating along independent trajectories, and a condition we do not assume.
@@
More recent methods [13, 30, 17] exploit high-fidelity 3D scene representations [19, 48] to model the pre-change scene and render novel views from post-change viewpoints, motivating pose-agnostic SCD. MV3DCD [13] showed that learning change information across multiple viewpoints with a scene representation significantly outperforms pairwise predictions. However, these approaches require a complete set of pre- and post-change captures, and use Structure-from-Motion [45] (SfM) to register poses to a common reference frame, confining them to an offline setting. In contrast, we study SCD in an online and incremental regime, where changes are inferred on-the-fly.
@@
Pose-agnostic anomaly detection [52, 22, 29] also builds a 3D representation of a pre-change object. To detect anomalies, the object is rendered from the post-change viewpoint, then scored using feature comparisons. However, these works focus on single objects rather than large-scale scenes. Approaches such as [22, 52] optimize camera poses directly against the representation, leading to slower pose estimation and, as shown by MV3DCD [13], frequent convergence failures in large complex scenes with multiple changes and view-dependent inconsistencies. Liu et al. [29] improved efficiency by replacing this step with HLoc [42].
While also employing a self-supervised objective, the approach of Furukawa et al. [12] differs fundamentally from our formulation of the self-supervised loss. Their loss is designed to exclude high-error regions to facilitate 2D alignment, whereas ours instead integrates complementary yet potentially noisy change cues across modalities into a persistent 3D representation.
@@
ChangeSim [34] also investigates online SCD. However, ChangeSim depends on an off-the-shelf RGB-D SLAM system [23] for pose estimation and assumes that pre- and post-change trajectories are closely aligned. This reduces the task to image retrieval, where the nearest pre-change view (by $L_{1}$ distance between camera poses) is selected. In contrast, we make no assumptions about incoming RGB-only frames or the trajectories; instead, we estimate poses directly in the pre-change coordinate frame and infer change masks by jointly leveraging all viewpoints observed so far. Our approach operates fully label-free, pose-agnostic, multi-view, online, and at real-time rates.
@@ 2.2 Efficient Representation Update
NeRFs [32] and 3DGS [19] are widely adopted photorealistic scene representations, capturing fine geometry and appearance. NeRFs regress a 5D plenoptic function [4] using an MLP network to parameterize density and view-dependent radiance, while 3DGS employs anisotropic Gaussian primitives for real-time novel-view-synthesis.
Recently, there has been growing interest in real-time reconstruction [31, 27]. However, these methods generally underperform compared to offline counterparts and require substantial view overlap between frames. In contrast, approaches that address the long-term evolution of scenes remain less explored, focusing on updating representations from sparse and intermittent captures [1, 49]. Closely related is continual learning for photorealistic scene representations [1, 51, 5, 49]. NeRF-based continual learning approaches [49, 5] utilize distillation or generative replay but inherit slow inference and longer optimization times. GaussianUpdate [51] proposes a three-stage optimization pipeline requiring substantial training iterations. CL-Splats [1] introduces a local optimization kernel to calculate gradients only for changed primitives, yet cannot robustly handle global appearance variations, particularly the illumination shifts often present between real-world inspection scenarios.
Long-term scene representations are also well-explored in robotics for autonomous navigation. To maintain spatial consistency in dynamic environments, these systems utilize volumetric representations [44, 11], object-aware tracking [36, 53], unified metric-semantic frameworks [43], and recently, 3DGS [50].
Continual learning approaches [1, 51, 49, 5] focus on updating representations while facilitating history recovery. In contrast, we focus on updating the representation with minimal training overhead, enabling frequent repeated inspections. To this end, we propose a simple selective modeling strategy that only reconstructs changed regions, guided by our change masks and the change representation, followed by fusion with existing primitives. A lightweight global optimization step ensures consistency, enabling updates within seconds while robustly handling both geometric and appearance changes, including global illumination variations.
@@ 3 Methodology
Our approach is illustrated in Fig. 3. We begin by constructing a 3DGS [19] representation of the pre-change (reference) scene offline (Sec. 3.1). Incoming images of the post-change (inference scene) are processed online. We estimate its pose relative to the reference scene (Sec. 3.2), then render the corresponding viewpoint to extract change cues (Sec. 3.3). These cues are used to infer a change mask (Sec. 3.4), leveraging current and previous observations. After processing all observations, the representation is updated (Sec. 3.5) to the current state of the environment.
Problem Setup: The reference scene $\mathcal{R}_{\text{ref}}$ is captured with a set of $n_{\text{ref}}$ images, $\mathcal{I}_{\text{ref}}=\{I_{\text{ref}}^{k}\}_{k=1}^{n_{\text{ref}}}$. Over time, the scene undergoes changes in structure (e.g., additions, removals, or object movement), or object appearance (e.g., variations in color or texture), forming the inference scene $\mathcal{R}_{\text{inf}}$. In addition, there may be ‘distractors’ such as reflections, shadows and global illumination changes.
@@ 3.1 Building Reference Scene Representation
Following the standard 3DGS pipeline [19], we first estimate the camera poses $\mathcal{P}_{\text{ref}}=\{P_{\text{ref}}^{k}\}_{k=1}^{n_{\text{ref}}}$ for all images in $\mathcal{I}_{\text{ref}}$ using SfM [45]. Using $\mathcal{I}_{\text{ref}}$, $\mathcal{P}_{\text{ref}}$ and a sparse point cloud from SfM, we construct the $\mathcal{R}_{\text{ref}}$ with Speedy-Splat [15]. We assume that the viewpoints, scene coverage, and image quality in $\mathcal{I}_{\text{ref}}$ are sufficient to produce a high-fidelity $\mathcal{R}_{\text{ref}}$.
@@ 3.2 Inferring Pose of an Incoming Frame
For each reference image $I_{\text{ref}}^{k}\in\mathcal{I}_{\text{ref}}$, we extract keypoints and descriptors using XFeat [35] as a fast, lightweight detector, followed by exhaustive matching across all reference images. Using the known camera poses $\mathcal{P}_{\text{ref}}$, these correspondences are triangulated to form a consistent 3D point set associated with each $I_{\text{ref}}^{k}$. This point set serves as a geometric anchor for estimating the pose of incoming inference frames by establishing 2D–3D correspondences between their detected keypoints and the reference points.
Given an incoming frame $I_{\text{inf}}^{k}$, we extract its descriptors and select the top-$n$ reference frames with the highest number of matches ($n=4$ in our experiments). These reference frames provide candidate 2D–3D correspondences, which are then used to estimate the pose $P_{\text{inf}}^{k}$ of $I_{\text{inf}}^{k}$ via PnP with RANSAC [26, 10]. Finally, we refine $P_{\text{inf}}^{k}$ with inliers through a GPU-parallel miniBA [31].
Since pose estimation for $I_{\text{inf}}^{k}$ relies exclusively on the retrieved reference frames, the system operates without drift accumulation. Moreover, by restricting inference to a fixed-size set of reference frames, we achieve constant-time $O(1)$ pose estimation. We discuss some limitations of the XFeat features in the supplementary material.